OpenStudy (anonymous):

Which ratio is equivalent to 12 : 32?

OpenStudy (anonymous):

Answers?

OpenStudy (anonymous):

is would be 3:8

OpenStudy (anonymous):

7/15 48/128 28/63 48/64

OpenStudy (solomonzelman):

\[12=3 \times 2 \times 2\]\[32=2 \times 2 \times 2 \times 2 \times 2\]

OpenStudy (anonymous):

B

OpenStudy (anonymous):

When you multiply both by 4 you get 48:128

OpenStudy (anonymous):

48/128

OpenStudy (anonymous):

Thank you!